# Asynchronous mass inversion enriched quantum anomalous Hall states in multilayer graphene

### v1

@@ Abstract
Recently, multilayer graphene systems have attracted significant attention due to the discovery of a variety of intriguing phases, particularly quantum anomalous Hall (QAH) states. In rhombohedral pentalayer graphene, both $C=-5$ and $C=-3$ QAH states have been observed. While the $C=-5$ QAH state is well understood, the origin of the $C=-3$ QAH state remains unclear. In this letter, we propose that the $C=-3$ QAH state in rhombohedral pentalayer graphene (RPG) arises from an asynchronous mass inversion mechanism, driven by the interplay between trigonal warping and staggered layer order under an applied displacement field. Trigonal warping splits the low-energy bands into a central touching point and three “leg” Dirac cones. In the presence of staggered layer order, this splitting enables mass inversions driven by the displacement field to occur asynchronously at the central touching point and the “leg” Dirac cones, potentially leading to the formation of the $C=-3$ QAH state. Furthermore, this mechanism can also be applied to Bernal multilayer graphene systems, predicting the existence of additional QAH states beyond $C=\pm N,\pm 2N$ for $N$-layer graphene.
Introduction.—Rhombohedral multilayer graphene (RMG) has become a focal point in condensed matter physics, attracting significant attention both theoretically and experimentally. It serves as an ideal platform for investigating a variety of intriguing phases, including unconventional superconductivity [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16], integer and fractional quantum anomalous Hall (QAH) states [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44], ferro-valleytricity order [45, 46], and so on. Recently, both $C=-5$ and $C=-3$ QAH states have been observed in rhombohedral pentalayer graphene (RPG) [21]. The emergence of the $C=-5$ QAH state can be understood through mass inversion in a single valley with a single spin, considering the presence of layer-antiferromagnetic (LAF) order and the displacement field in a generalized Dirac model with $k^{5}$ dispersion [21, 25, 47, 48]. However, the origin of the $C=-3$ QAH state remains a puzzle until now.
In this letter, we propose an asynchronous mass inversion mechanism to explain the emergence of $C=-3$ QAH state in RPG. As discussed in previous works [49, 50, 51, 17, 52, 53, 47], the low-energy bands near two valleys ($K$ and $K^{\prime}$) for RPG are primarily contributed by the B sublattice of the top layer and the A sublattice of the bottom layer, both of which are non-dimerized. The side view and hopping parameters for RPG are shown in Fig. 1 (a). In the presence of trigonal warping, the low-energy bands of RPG near the $K$ (or $K^{\prime}$) valley split into three “leg” Dirac cones with $C_{3}$ symmetry, along with a central point exhibiting quadratic band-touching, as shown in Fig. 1 (b). Additionally, by analogy to the LAF order [54, 51, 55, 56, 57, 58, 59, 60, 61, 62], we introduce a staggered layer order [63], where in even-layer graphene, the charge densities between adjacent layers are oppositely polarized, and in odd-layer graphene, the middle layer lacks an order parameter, while adjacent layers on either side of the middle layer exhibit oppositely polarized charge densities. This order can be induced by Coulomb interaction near charge neutrality [51, 64, 61, 65, 66, 67, 68, 69, 70]. The coexistence of staggered layer order and the displacement field results in a momentum-dependent effective mass in the low-energy two-band model for RPG. With a fixed amplitude of staggered layer order, the central band-touching point and the “leg” Dirac cones, having different momenta, will exhibit different effective masses under the same displacement field. Consequently, mass inversion for these two types of band-touching points will occur asynchronously at different displacement fields. When mass inversion occurs at the three “leg” Dirac cones, the change in the Chern number of RPG can be ±$\pm 3$, as each “leg” Dirac cone carries a Berry phase of $\zeta\pi$ ($\zeta=\pm 1$, corresponding to the $K$ and $K^{\prime}$ valleys). This mechanism may give rise to a QAH state with $C=-3$ in RPG. As the displacement field increases further, mass inversion occurs at the central touching point. Due to the quadratic band-touching, this point has a Berry phase of $2\zeta\pi$. Therefore, mass inversion at this point results in a change in the Chern number of ±$\pm 2$, which could drive a topological phase transition from the QAH state with $C=-3$ to $C=-5$. Furthermore, this mechanism can also be applied to the Bernal multilayer graphene (BMG) system. Our theory solves the long-standing problem of $C=-3$ QAH state in RPG, and provides a minimal model to understand the $C\neq\pm N,\pm 2N$ QAH states in $N$-layer graphene for future experiments.
@@
Figure 1: (a). Side view of the rhombohedral pentalayer graphene (RPG), the blue and red atoms represent the $A$ and $B$ sublattices, respectively. The hopping parameters are labeled by different colors ($v_{i}=\sqrt{3}\gamma_{i}/2$): $\gamma_{0}=3160~{}meV$, $\gamma_{1}=390~{}meV$, $\gamma_{2}=-20~{}meV$, $\gamma_{3}=315~{}meV$, and $\gamma_{4}=44~{}meV$. (b). The band structure near $K$ (or $K^{\prime}$) valley without (left) and with (right) trigonal warping. The presence of trigonal warping leads to the low-energy bands of RPG split into three “leg” Dirac cones and a central point with quadratic band-touching. (c). The band structure of the low-energy two-band model in Eq. (2) along $k_{y}=0$ line at $K$ valley. The red solid lines represent the bands for the case where $m=0$, $\Delta=0$, and $\gamma_{2}=-20~{}meV$, while the blue solid lines represent the bands in the case where $m=-10~{}meV$, $\Delta=5~{}meV$ and $\gamma_{2}=-20~{}meV$. (d). The band structure of the low-energy two-band model in Eq. (2) along $k_{y}=0$ line at $K$ valley when the trigonal warping effect is absent ($\gamma_{2}=0$). In (c) and (d), the same color of the bands indicates the same values of $m$ and $\Delta$.
The low-energy two-band model for RPG—We start by introducing the low-energy two-band effective model for RPG in the presence of staggered layer order and a displacement field. The specific form of the Hamiltonian for the staggered layer order, expressed in the basis $\psi=(A_{1},B_{1},A_{2},B_{2},\dots,A_{5},B_{5})$, is given by:
@@
where $U_{l}=(-1)^{l-1}\sigma_{0}$, with $\sigma_{0}$ being the two-dimensional identity matrix. $l$ is the layer index, and $O$ is the two-dimensional zero matrix. $Diag(...)$ denotes a diagonal matrix whose diagonal elements are listed in the brackets, and $m$ is the amplitude of this order.
By performing perturbation theory, the Hamiltonian of the low-energy two-band model for RPG under the basis $(A_{1},B_{5})$ is written as (see Supplemental Material for details):
@@
where $X_{5}(\mathbf{k})$ is the effective hopping between $A_{1}$ and $B_{5}$, its specific form is written as:
@@
In this equation, $\xi_{\mathbf{k}}=\zeta k_{x}+ik_{y}$ ($\zeta=\pm$ for $K$ and $K^{\prime}$ valley), $a_{0}$ is the lattice constant of graphene, $v_{0}$ represents the intralayer hopping, $\gamma_{1}$ represents the hopping between dimerized atoms belonging to adjacent layers, and $\gamma_{2}$ represents the next-nearest interlayer hopping. The band structures of the low-energy effective model with ($\gamma_{2}\neq 0$) and without ($\gamma_{2}=0$) the trigonal warping effect along $k_{y}=0$ line are shown in Fig 1 (c) and (d), respectively.
Additionally, the effective mass $M_{5,eff}(k)$ of the two-band model for RPG is given by:
$M_{5,eff}(k)=(m+\Delta)-\frac{(2m-\Delta)v_{0}^{2}a_{0}^{2}k^{2}}{2\gamma_{1}^{2}},$ (4)
@@
where $m$ is the amplitude of the staggered layer order, $\Delta$ is the interlayer potential induced by the displacement field, and $k$ is the amplitude of the momentum measured from the $K$ (or $K^{\prime}$) point of the Brillouin zone (BZ).
@@
Quantum anomalous Hall states in RPG—The effective mass given in Eq. (4) is momentum-dependent. At the central point of $K$ or $K^{\prime}$ valley, the effective mass is expressed as $m+\Delta$. It changes sign when the displacement field satisfies $\Delta=\Delta_{1}=-m$ (assuming $m<0$). At the point with $k\neq 0$, the effective mass changes sign at $\Delta=\Delta_{2}=\frac{2\lambda-1}{1+\lambda}m$, where $\lambda=\frac{v_{0}^{2}a_{0}^{2}k^{2}}{2\gamma_{1}^{2}}$. When $\lambda$ satisfies the condition $\lambda<\frac{1}{2}$, it follows that $0<\Delta_{2}<\Delta_{1}$. As the displacement field increases, the effective mass of the three “leg” Dirac cones changes sign first at $\Delta=\Delta_{2}$, followed by the sign change of the effective mass at the central quadratic touching point at $\Delta=\Delta_{1}$. The evolution of the band structure with increasing interlayer potential for $K^{\prime}$ valley is shown in Fig. 2 (a). To explain the appearance of the QAH state with $C=-3$, we consider staggered layer orders with amplitudes $m_{1}$, $-m_{1}$, $m_{2}$, and $-m_{2}$ for the four flavors $K\uparrow$, $K\downarrow$, $K^{\prime}\uparrow$, and $K^{\prime}\downarrow$, respectively, where $m_{1}\ll m_{2}<0$. When the interlayer potential satisfies $\Delta=\Delta_{2}$, a mass inversion occurs at the ”leg” Dirac cones, causing the Chern number of the valence band with spin-up in the $K^{\prime}$ valley (flavor $K^{\prime}\uparrow$) to change from $\frac{5}{2}$ to $-\frac{1}{2}$. Meanwhile, the Chern numbers for the valence band with the other three flavors—$K\uparrow$, $K\downarrow$, and $K^{\prime}\downarrow$—are $-\frac{5}{2}$, $\frac{5}{2}$, and $-\frac{5}{2}$, respectively, and remain unchanged. Thus, the QAH state with $C=-3$ appears. When the interlayer potential further increases beyond $\Delta_{1}=-m_{2}$, the effective mass at the central band-touching point for the $K^{\prime}\uparrow$ flavor changes, leading to the emergence of a QAH state with $C=-5$.
@@
Figure 2: (a). The evolution of the band structure for the low-energy two-band model Eq. (2) along $k_{y}=0$ line with increasing interlayer potential at $K^{\prime}$ valley, in the presence of staggered layer order Eq. (1), is shown, where the parameters are $v_{0}=\frac{\sqrt{3}}{2}\times 3160~{}meV$, $\gamma_{1}=390~{}meV$, $m=m_{2}=-10~{}meV$ and $\gamma_{2}=-20~{}meV$. (b). The evolution of the band structure with increasing interlayer potential in the case where $\gamma_{2}=0$ (without trigonal warping). All other parameters are the same as in (a). (c). The evolution of the band structure with increasing interlayer potential at $K$ valley, in the presence of the staggered sublattice order given in Eq. (5). The parameters are $m=m_{1}=-15~{}meV$ and $\gamma_{2}=-20~{}meV$. (d). The evolution of the band structure with increasing interlayer potential in the case where $\gamma_{2}=0$. All other parameters are the same as in (c).
@@
For contrast, we also show the evolution of the band structure with increasing interlayer potential in the case where $\gamma_{2}=0$ (without trigonal warping) in Fig. 2 (b). As the interlayer potential increases, the effective mass at the central point of $K^{\prime}$ valley will change sign at $\Delta=\Delta_{1}=-m_{2}$, with no other gap-closing occurring. Additionally, the dispersion of the band-touching point follows the form $k^{5}$. Therefore, in the absence of trigonal warping, increasing the interlayer potential induces a single topological phase transition from the $C=0$ state to a QAH state with $C=-5$. In addition, if there is no staggered layer order ($m=0$), although this effective mass (Eq. 4) remains momentum-dependent, its sign is only determined by the sign of the interlayer potential $\Delta$ at both $k=0$ and $k\neq 0$ points. Therefore, increasing the interlayer potential (from negative to positive) will induce mass inversion at both the central point and the “leg” Dirac cones simultaneously.
The entire phase diagram under increasing displacement field for RPG—Using the staggered layer order defined in Eq. (1), we can explain the topological phase transitions from the $C=0$ state to the $C=-3$ QAH state, and then to the $C=-5$ QAH state. However, as the displacement field continues to increase, the transition from the $C=-5$ to the $C=-3$ QAH state, and then to the $C=0$ state, remains unexplained. To account for this sequence of topological phase transitions, it is necessary to introduce an order parameter that first induces mass inversion at the central point. One natural choice is the staggered sublattice order, which describes a uniform charge density across each layer but with opposite polarizations between the $A$ and $B$ sublattices within a single layer. This order also satisfies the charge neutrality condition. The specific form of the Hamiltonian for this order parameter in the basis $\psi=(A_{1},B_{1},A_{2},B_{2},\dots,A_{5},B_{5})$ is written as:
$H^{\prime}_{m}=m~{}Diag(\sigma_{z},\sigma_{z},...,\sigma_{z}),$ (5)
where $\sigma_{z}$ is the Pauli matrix of the sublattice degree of freedom on each layer. By performing the perturbation theory, the effective mass for the low-energy effective model is derived as:
$M^{\prime}_{5,eff}(k)=(m+\Delta)+\frac{(2m+\Delta)v_{0}^{2}k^{2}}{2\gamma_{1}^{2}}.$ (6)
The hopping term for the low-energy two-band model is the same as Eq.(3). For $k=0$ point, the effective mass changes sign at $\Delta=\Delta^{\prime}_{1}=-m$ (assuming $\Delta>0$ and $m<0$). However, for $k\neq 0$ points, unlike the case of the staggered layer order in Eq. (1), the effective mass changes sign at $\Delta=\Delta^{\prime}_{2}=-m\frac{1+2\lambda}{1+\lambda}$ which is larger than $\Delta^{\prime}_{1}=-m$ due to $\lambda=\frac{v_{0}^{2}a_{0}^{2}k^{2}}{2\gamma_{1}^{2}}>0$. Thus, as the displacement field increases, the effective mass at the central point of the $K$ (or $K^{\prime}$) valley changes sign first at $\Delta=\Delta_{1}^{\prime}=-m$, followed by the sign change of the effective mass of the three “leg” Dirac cones at $\Delta=\Delta^{\prime}_{2}>\Delta_{1}^{\prime}$. In the presence of the staggered sublattice order, the evolution of the band structure at the $K$ valley with increasing interlayer potential is shown in Fig. 2 (c) and (d), for cases with and without trigonal warping, respectively.
We then assume that the staggered orders for the four flavors $K\uparrow$, $K\downarrow$, $K^{{}^{\prime}}\uparrow$, and $K^{{}^{\prime}}\downarrow$ are given as $H^{\prime}_{m_{1}}$, $-H^{\prime}_{m_{1}}$, $H_{m_{2}}$, and $-H_{m_{2}}$, respectively, with $0>m_{2}>m_{1}$. When the interlayer potential satisfies $0<\Delta<-m_{2}$, as $\Delta$ increases, a mass inversion happens at the three “leg” Dirac cones for $K^{\prime}\uparrow$ flavor, causing the Chern number of RPG to change from $C=0$ to $C=-3$. Then, when $\Delta=-m_{2}$, the effective mass at the central band-touching point for electrons with flavor $K^{\prime}\uparrow$ changes sign, leading to a further change in the Chern number from $C=-3$ to $C=-5$. When the displacement field satisfies $-m_{2}<\Delta<-m_{1}$, since $m_{2}>m_{1}$, the Chern number for RPG remains $C=-5$. As the displacement field continues to increase, for electrons with flavor $K\uparrow$, the effective mass of the central band-touching point changes sign at $\Delta=-m_{1}$, leading to a topological phase transition from the QAH state with $C=-5$ to $C=-3$. With further increases in the displacement field, the effective masses of the three “leg” Dirac cones with flavor $K\uparrow$ change sign, resulting in a topological phase transition from the QAH state with $C=-3$ to the state with $C=0$. Consequently, we deduce that as the displacement field increases, the Chern number for RPG evolves as $0\rightarrow-3\rightarrow-5\rightarrow-3\rightarrow 0$, which is consistent with experimental observations [21]. Moreover, within this mechanism, we can deduce that at $\Delta=0$, the $C=0$ state in RPG corresponds to a state where both the spin and valley Chern numbers are 0. When the displacement field becomes sufficiently large, although the total Chern number of RPG remains 0, the system transitions into a quantum valley Hall state with a valley Chern number of 5. Additionally, to make the analysis more practical, we investigate the band structure and topological phase transition using the full Hamiltonian of RPG, taking into account the staggered layer (or sublattice) order and displacement field. The asynchronous mass inversion at the central touching point and ”leg” Dirac cones remains unchanged (see Supplemental Material for details). Therefore, the low-energy effective model in Eq. (2) captures the key physical mechanism underlying the QAH states in RPG.
The low-energy effective model for BMG—Next, we apply this asynchronous mass inversion mechanism to explore the possible enriched QAH states in BMG. For $N$-layer BMG, similar to the case of RMG, the combination of trigonal warping and staggered layer order also gives rise to additional QAH states with Chern number $C\neq\pm N,\pm 2N$ when a displacement field is applied. However, unlike the case in RMG, the low-energy effective model for BMG is no longer a two-band model. For Bernal $N$-layer graphene, if $N$ is even, the Hamiltonian can be decomposed into $N/2$ mutually coupled bilayer graphene systems [71]. We can then project the Hamiltonian for each bilayer graphene onto the basis which is composed of non-dimerized “atoms” to obtain an $N$-dimensional low-energy effective model (see Supplemental Material for details). If $N$ is odd, the low-energy effective model for BMG consists of $N/2$ mutually coupled bilayer graphene systems along with an additional monolayer graphene that couples to these bilayer graphenes. The presence of this additional monolayer graphene results in a metallic state, even when staggered layer orders are considered [63], thereby preventing the formation of a QAH state. As a result, this paper focuses on even-layer BMGs. Specifically, we use Bernal tetralayer graphene (BTG) as an example in the following discussion. The side view and hopping parameters for BTG are shown in Fig. 3 (a),

### v2

@@ Abstract
Recently, multilayer graphene systems have attracted significant attention due to the discovery of a variety of intriguing phases, particularly quantum anomalous Hall (QAH) states. In rhombohedral pentalayer graphene (RPG), both QAH states with Chern number $C=-5$ and $C=-3$ have been observed. While the $C=-5$ QAH state is well understood, the origin of the $C=-3$ QAH state remains unclear. In this letter, we propose that the $C=-3$ QAH state, as well as the topological phase transition from $C=-3$ to $C=-5$ state in RPG, arises from an asynchronous mass inversion mechanism driven by the interplay between trigonal warping, staggered layer order, and the displacement field: Trigonal warping splits the low-energy bands of RPG into a central touching point and three satellite Dirac cones. Meanwhile, the coexistence of the staggered layer order and displacement field induces a momentum-dependent effective mass in the low-energy bands. Consequently, mass inversions at the central touching point and the satellite Dirac cones, induced by an increasing displacement field, can occur asynchronously, leading to the formation of the $C=-3$ QAH state and the topological phase transition from QAH state with $C=-3$ to $C=-5$. Additionally, based on this mechanism, we predict the presence of a $C=3$ QAH state in rhombohedral tetralayer graphene (RTG), which can be detected experimentally. Furthermore, this mechanism can also be applied to Bernal tetralayer graphene (BTG), explaining the origin of the observed $C=6$ QAH state.
Introduction.—In recent years, multilayer graphene has been widely studied both theoretically and experimentally, since it provides an ideal platform for exploring several intriguing phases, such as the unconventional superconducting phase Zhou et al. (2021); Ghazaryan et al. (2021); You and Vishwanath (2022); Zhou et al. (2022); Chou et al. (2022); Pangburn et al. (2023); Qin et al. (2023); Wagner et al. (2024); Li et al. (2024); Yang et al. (2024); Viñas Boström et al. (2024); Fischer et al. (2024); Geier et al. (2024); Chau et al. (2024); Wang et al. (2024a, b); Kim et al. (2025); Choi et al. (2025); Zhang and Wang (2025); Han et al. (2025), the integer and fractional quantum anomalous Hall (QAH) states Nandkishore and Levitov (2010); Zhang et al. (2010a, 2011); Shi et al. (2020); Kerelsky et al. (2021); Park et al. (2023); Kwan et al. (2023); Mondal et al. (2023); Yu et al. (2024); Liu et al. (2024a); Dong et al. (2024a); Han et al. (2024a, b); Sha et al. (2024); Liu and Wang (2025); Dong et al. (2024b); Soejima et al. (2024); Dong et al. (2024c); Kudo et al. (2024); Perea-Causin et al. (2024); Xie and Das Sarma (2024); Lu et al. (2024a); Ju et al. (2024); Huang et al. (2024a); Das Sarma and Xie (2024); Herzog-Arbeitman et al. (2024); Xie et al. (2024); Huang et al. (2024b); An et al. (2024); Lu et al. (2024b); Zhou et al. (2024); Aronson et al. (2024); Kwan et al. (2024); Wei et al. (2025); Lu et al. (2025); Pichler et al. (2025); Waters et al. (2025); Bernevig and Kwan (2025), and the ferro-valleytricity ordered state Han et al. (2023); Liu et al. (2024b); Islam et al. (2025). Among these fascinating phases, QAH states with high Chern numbers in rhombohedral multilayer graphene (RMG) are of particular interest due to their potential applications in tunable low-dissipation electronic devices Han et al. (2024a, b); Sha et al. (2024); Liu and Wang (2025). The possibility of realizing QAH states in RMG was proposed over a decade ago Zhang et al. (2011), based on a mechanism extended from the Haldane model Haldane (1988), where the inversion of the Dirac mass at a single valley plays a central role Mele (2015); Weng et al. (2015); Liu et al. (2016); Chang et al. (2023). In rhombohedral $N$-layer graphene, the low-energy physics is described by a two-band Dirac model with a $|k|^{N}$ dispersion McCann and Fal’ko (2006); Min and MacDonald (2008); Nakamura and Hirasawa (2008); Grüneis et al. (2008); Koshino and McCann (2009); Zhang et al. (2010b); Jia et al. (2013); McCann and Koshino (2013), where the relevant orbital components predominantly originate from the bottom-layer $A_{1}$ sublattice and the top-layer $B_{N}$ sublattice. Within this two-band model, the presence of layer-antiferromagnetic (LAF) order Nilsson et al. (2006); Zhang et al. (2010a); Lemonik et al. (2012); Lang et al. (2012); Scherer et al. (2012); Kharitonov (2012); Wang et al. (2013); Yuan et al. (2013); Liang et al. (2022); Jiang et al. (2024) and Ising spin-orbit coupling (SOC) Gmitra et al. (2013); Gmitra and Fabian (2015, 2017); Kochan et al. (2017); Li and Koshino (2019); Naimer et al. (2021); Zollner et al. (2023) enables the displacement field to drive a mass inversion in the band associated with a single spin and valley. This leads to the emergence of QAH states with Chern number $C=\pm N$ Zhang et al. (2011); Liu and Wang (2025).
Recently, the experimental realization of the QAH state in rhombohedral pentalayer graphene (RPG) Han et al. (2024a) has confirmed the predicted QAH state with Chern number $C=-5$. Surprisingly, another QAH state with $C=-3$ has also been observed in RPG, which remains unexplained by the current theoretical framework Zhang et al. (2011); Jung et al. (2011); Han et al. (2024a); Liu and Wang (2025). In this Letter, we identify the crucial role of the trigonal warping effect, which is usually neglected when investigating the QAH state in multilayer graphene systems Zhang et al. (2011); Jung et al. (2011); Han et al. (2024a); Liu and Wang (2025). By incorporating trigonal warping into the low-energy effective Hamiltonian of RPG, we propose that the $C=-3$ QAH state observed in RPG arises from an asynchronous mass inversion mechanism, described as follows: In the presence of trigonal warping, the low-energy bands near the $K$ (or $K^{\prime}$) valley split into three satellite Dirac cones with $C_{3}$ symmetry, along with a central quadratic band touching point Koshino and McCann (2009), as illustrated in Fig.1(a). When a staggered layer order Grushina et al. (2015) (which is analogous to the layer antiferromagnetic (LAF) order) coexists with the displacement field, their combined effect induces a momentum-dependent effective mass in the low-energy bands of RPG. As a result, under a fixed staggered layer order, the central band-touching point and the satellite Dirac cones acquire different effective masses under the same displacement field, owing to their distinct momenta. Consequently, as the displacement field increases, mass inversions at the central and satellite band-touching points occur asynchronously. When mass inversion occurs at the three satellite Dirac cones, because each of them carries a Berry phase of $\zeta\pi$ (with $\zeta=\pm 1$ for the $K$ and $K^{\prime}$ valleys, respectively), the Chern number in RPG can change by ±$\pm 3$. This mechanism leads to the emergence of a QAH state with $C=-3$ in RPG.
Additionally, the asynchronous mass inversion mechanism also accounts for the topological phase transition from the $C=-3$ to $C=-5$ QAH state observed in RPG as the displacement field increases Han et al. (2024a). After the $C=-3$ QAH state appears, a further increase in the displacement field leads to a mass inversion at the central band-touching point. Owing to its quadratic nature, the band near this point carries a Berry phase of $2\zeta\pi$. As a result, its mass inversion changes the Chern number by ±$\pm 2$, driving a topological phase transition from the $C=-3$ QAH state to $C=-5$. Importantly, the asynchronous mass inversion mechanism is generally applicable to multilayer graphene systems, regardless of stacking type (rhombohedral or Bernal) or layer number. For example, our theory also explains the observed $C=6$ QAH state Cho et al. (2025) in Bernal-stacked tetralayer graphene (BTG). Our theory addresses the long-standing puzzle regarding the origin of the $C=-3$ QAH state in RPG, and offers a possible mechanism for understanding $C\neq\pm N,\pm 2N$ QAH states in $N$-layer graphene, thereby guiding future experimental explorations.
@@
Figure 1: (a). The band structure for the low-energy two-band model near $K$ (or $K^{\prime}$) valley without (left) and with (right) trigonal warping. The presence of trigonal warping leads to the low-energy bands of RPG split into three satellite Dirac cones and a central point with quadratic band-touching. (b). Side view of the rhombohedral pentalayer graphene (RPG), the blue and red atoms represent the $A$ and $B$ sublattices, respectively. The hopping parameters are labeled by different colors ($v_{i}=\sqrt{3}\gamma_{i}/2$): $\gamma_{0}=3160~{}meV$, $\gamma_{1}=390~{}meV$, $\gamma_{2}=-20~{}meV$, $\gamma_{3}=315~{}meV$, and $\gamma_{4}=44~{}meV$ Koshino and McCann (2009). (c). The band structure of the low-energy two-band model in Eq. (4) along $k_{y}=0$ line at $K$ valley. The red solid lines represent the bands for the case where $m=0$, $\Delta=0$, and $\gamma_{2}=-20~{}meV$, while the blue solid lines represent the bands in the case where $m=-10~{}meV$, $\Delta=5~{}meV$ and $\gamma_{2}=-20~{}meV$. (d). The band structure of the low-energy two-band model in Eq. (4) along $k_{y}=0$ line at $K$ valley when the trigonal warping effect is absent ($\gamma_{2}=0$). In (c) and (d), the same color of the bands indicates the same values of $m$ and $\Delta$.
The effective model for RPG— The Hamiltonian for RPG without the displacement field and orders in the basis $\psi=(A_{1},B_{1},A_{2},B_{2},\dots,A_{5},B_{5})$ can be expressed as:
$H_{RPG}=\begin{pmatrix}H_{0}&V&W&&\\ V^{\dagger}&H_{0}&V&W&\\ W^{\dagger}&V^{\dagger}&H_{0}&V&W\\ &W^{\dagger}&V^{\dagger}&H_{0}&V\\ &&W^{\dagger}&V^{\dagger}&H_{0}\end{pmatrix},$ (1)
where $A_{i}$ and $B_{i}$ represent the $A$ and $B$ sites at layer $i$. Here, $H_{0}$ is a $2\times 2$ matrix that represents the intralayer hopping, while $V$ and $W$ encode the nearest and next-nearest interlayer hopping terms, respectively. Their specific form can be found in the Supplemental Material sup , and the hopping terms are shown in Fig. 1(b). The interlayer potential induced by the displacement field is written as Ghazaryan et al. (2023):
$H_{\Delta}=\Delta~{}Diag(D_{1},D_{2},...,D_{5}),$ (2)
where $Diag(...)$ denotes a diagonal matrix whose diagonal elements are listed in the brackets, $D_{l}=(1-\frac{(l-1)}{2})\sigma_{0}$ (for $l=1,2,...,5$), $\sigma_{0}$ is the $2\times 2$ identity matrix, $l$ is the layer index, and $\Delta$ is the amplitude of this interlayer potential. Additionally, the staggered interlayer order is:
@@
where $U_{l}=(-1)^{l-1}\sigma_{0}$, $O$ is the $2\times 2$ zero matrix, and $m$ is the amplitude of this order. To capture the key physical mechanism underlying the QAH states in RPG, we derive a low-energy two-band model from the full Hamiltonian ($H^{full}=H_{RPG}+H_{\Delta}+H_{m}$), which incorporates the displacement field and staggered layer order. Under the basis $(A_{1},B_{5})$, the low-energy model is expressed as:
@@
where $X_{5}(\mathbf{k})$ is the effective hopping between $A_{1}$ and $B_{5}$. The specific form of $X_{5}(\mathbf{k})$ is written as:
@@
In Eq. (5), $\xi_{\mathbf{k}}=\zeta k_{x}+ik_{y}$ ($\zeta=\pm$ for $K$ and $K^{\prime}$ valley), $a_{0}$ is the lattice constant of graphene, $v_{0}$ represents the intralayer hopping, $\gamma_{1}$ represents the hopping between dimerized atoms belonging to adjacent layers, and $\gamma_{2}$ represents the next-nearest interlayer hopping which induces the trigonal warping effect. The band structures of the low-energy effective model with ($\gamma_{2}\neq 0$) and without ($\gamma_{2}=0$) the trigonal warping effect along $k_{y}=0$ line are shown in Fig 1 (c) and (d), respectively.
Additionally, the effective mass $M_{5,\text{eff}}(k)$ of the two-band model for RPG is given by:
$M_{5,\text{eff}}(k)=(m+\Delta)-\frac{(2m-\Delta)v_{0}^{2}a_{0}^{2}k^{2}}{2\gamma_{1}^{2}},$ (6)
@@
where $m$ is the amplitude of the staggered layer order, $\Delta$ is the interlayer potential induced by the displacement field, and $k$ is the amplitude of the momentum measured from the $K$ (or $K^{\prime}$) point of the Brillouin zone (BZ). As we will show below, the presence of the trigonal warping term $\gamma_{2}$ and the momentum-dependent effective mass account for the asynchronous mass inversion mechanism.
@@
Quantum anomalous Hall states in RPG—The effective mass given in Eq. (6) is momentum-dependent. At the central point of $K$ or $K^{\prime}$ valley, the effective mass is expressed as $m+\Delta$. It changes sign when the interlayer potential satisfies $\Delta=\Delta_{1}=-m$ (assuming $m<0$). At the point with $k\neq 0$, the effective mass changes sign at $\Delta=\Delta_{2}=\frac{2\epsilon-1}{1+\epsilon}m$, where $\epsilon=\frac{v_{0}^{2}a_{0}^{2}k^{2}}{2\gamma_{1}^{2}}$. When $\epsilon$ satisfies the condition $\epsilon<\frac{1}{2}$, it follows that $0<\Delta_{2}<\Delta_{1}$. Therefore, as the displacement field increases, the effective masses of the three satellite Dirac cones change sign first at $\Delta=\Delta_{2}$, followed by the sign change of the effective mass at the central quadratic touching point at $\Delta=\Delta_{1}$. The evolution of the band structure with increasing interlayer potential ($\Delta$) for $K^{\prime}$ valley is shown in Fig. 2 (a). To explain the appearance of the QAH state with $C=-3$, we consider staggered layer orders with amplitudes $m_{1}$, $-m_{1}$, $m_{2}$, and $-m_{2}$ for the four flavors $K\uparrow$, $K\downarrow$, $K^{\prime}\uparrow$, and $K^{\prime}\downarrow$, respectively, where $m_{1}<m_{2}<0$. When the interlayer potential satisfies $\Delta=\Delta_{2}=\frac{2\epsilon-1}{1+\epsilon}m_{2}$, a mass inversion occurs at the satellite Dirac cones, causing the Chern number of the valence band with spin-up in the $K^{\prime}$ valley (flavor $K^{\prime}\uparrow$) to change from $\frac{5}{2}$ to $-\frac{1}{2}$. Meanwhile, the Chern numbers for the valence band with the other three flavors—$K\uparrow$, $K\downarrow$, and $K^{\prime}\downarrow$—are $-\frac{5}{2}$, $\frac{5}{2}$, and $-\frac{5}{2}$, respectively, and remain unchanged. Thus, the QAH state with $C=-3$ appears. When the interlayer potential further increases beyond $\Delta_{1}=-m_{2}$, the effective mass at the central band-touching point for the $K^{\prime}\uparrow$ flavor changes sign, causing the Chern number of its valence band to become $-\frac{5}{2}$. As a result, a QAH state with $C=-5$ appears.
@@
Figure 2: (a). The evolution of the band structure for the low-energy two-band model Eq. (4) along $k_{y}=0$ line with increasing interlayer potential at $K^{\prime}$ valley, in the presence of staggered layer order Eq. (3), is shown, where the parameters are $v_{0}=\frac{\sqrt{3}}{2}\times 3160~{}meV$, $\gamma_{1}=390~{}meV$, $m=m_{2}=-10~{}meV$ and $\gamma_{2}=-20~{}meV$. (b). The evolution of the band structure with increasing interlayer potential in the case where $\gamma_{2}=0$ (without trigonal warping). All other parameters are the same as in (a).
@@
For comparison, we also present the evolution of the band structure with increasing interlayer potential in the case where $\gamma_{2}=0$, as shown in Fig. 2(b). As the interlayer potential increases, the effective mass at the central point of the $K^{\prime}$ valley changes sign at $\Delta=\Delta_{1}=-m$, without any additional gap closing. Furthermore, since the band-touching at this point follows a $k^{5}$ dispersion, it carries a Berry phase of $5\pi$. Therefore, in the absence of trigonal warping, increasing the interlayer potential induces a single topological phase transition from the $C=0$ state to a QAH state with $C=-5$. In addition, if there is no staggered layer order ($m=0$), although this effective mass (Eq. 6) remains momentum-dependent, its sign is only determined by the sign of the interlayer potential $\Delta$ at both $k=0$ and $k\neq 0$ points. Therefore, increasing the interlayer potential (from negative to positive) will simultaneously induce mass inversion at both the central point and the satellite Dirac cones.